Q1T.7 In a specific non-dispersive medium, a sinusoidal wave's angular frequency is proportional to its wavenumber. True or False?

Can you please explain why.
Answer #1

as we know

f = c/ L

1/L = f/c

also wave number

k = 2pi (1/ L)

k = 2pi * f / c

k = ( 2 pi / c) f...(i)

now angular frequency

w = 2 pi f

f = w/ (2pi)

putting in (i)

k = (1/c) w

so, wave number(k) is directly proportional to angular frequency (w)

========

Comment before rate in case any doubt, will reply for sure.. goodluck